### v3

# George Barmpalias State Key Lab of Computer Science

Abstract
A major open problem in computational complexity is the existence of a one-way function, namely a function from strings to strings which is computationally easy to compute but hard to invert. Levin (2023) formulated the notion of one-way functions from reals (infinite bit-sequences) to reals in terms of computability, and asked whether partial computable one-way functions exist. We give a strong positive answer using the hardness of the halting problem and exhibiting a total computable one-way function.
1 Introduction
A function is one-way if it is computationally easy to compute but hard to invert, even probabilistically. Many aspects of computer science such as computational complexity, pseudorandom generators and digital signatures rely on one-way functions from strings (finite binary sequences) to strings [20, 19]. Although their existence is not known, much research effort has focused on their implications and relations with other fundamental problems in computation [15, 14, 11]. Recent research has established strong connections between one-way functions and Kolmogorov complexity [26, 21, 22, 13].
Levin [20, §2.2] asked for the existence of a one-way function from reals (infinite binary sequences) to reals, which he defined in terms of probabilistic computability with respect to the uniform measure $\mu$ [^1] This was also discussed in [6]. These are partial computable functions $f$ which preserve algorithmic randomness and
the probability that $M$ inverts $f$, namely $f(M(y))=y$, is zero
for each probabilistic Turing machine $M$. We give a positive answer:
Theorem.
There exists a total computable one-way surjection $f$.
The hardness of inverting $f$ is based on the hardness of the halting problem $\emptyset^{\prime}$. As a result, every randomized continuous inversion of $f$ computes $\emptyset^{\prime}$.
Gács [9] independently constructed $f$ which is ‘one-way’ with respect to the domain instead of the range of $f$: for each partial computable $g$,
$\mu\big{(}{\{\hskip 0.85355pt{(x,r)}:{f(g(f(x),r))=f(x)}\hskip 0.85355pt\}}\big{)}=0$ (1)
and $f$ is partial computable with domain of positive uniform measure $\mu$.
Although (1) appears to be a closer analogue to the one-way functions in computational complexity, it does not preserve randomness, which is required by Levin [20]. Effective partial maps that meet (1) often have a null image. Levin’s definition implies (1) while the converse fails [5, §3.2].
In §3 we justify Levin’s definition as the correct analogue of one-way functions on discrete domains, over weaker formalizations. We exhibit a one-way real function using a framework that can be adapted toward meeting additional conditions, and explore their properties.
We do not know if there is an injective one-way real function. In §4 we explore the extent to which injectivity is compatible with one-way real functions and end with a summary and some problems in §5.
2 Preliminaries
Let $\mathbb{N}$ be the set of natural numbers, represented by $n,m,i,j,t,s$. Let
  • $2^{\omega}$ be the set of reals, represented by variables $x,y,z,v,w$
  • $2^{<\omega}$ the set of strings, represented by variables $\sigma,\tau,\rho$.
We index the bits $x(i)$ of $x$ starting from $i=0$. Let
  • $x\upharpoonright_{n}$ be the $n$-bit prefix $x(0)x(1)\cdots x(n-1)$ of $x$
  • $\preceq$, $\prec$ be the prefix and strict prefix relation between strings
  • $\sigma\mid\tau$ denotes that $\sigma\not\preceq\tau\ \wedge\ \tau\not\preceq\sigma$
and $\succeq$,$\succ$ , ≻ denote the suffix relations which, along with the prefix relations can also apply between a string and a real. Given $x,y$ let
$x\oplus y:=z$ where $z(2n)=x(n)$ and $z(2n+1)=y(n)$
and similarly for strings of the same length.
2.1 Computability and randomness
The Cantor space is $2^{\omega}$ with the topology generated by the basic open sets
$\llbracket\sigma\rrbracket:=\{\hskip 0.85355pt{z\in 2^{\omega}}:{\sigma\prec z}\hskip 0.85355pt\}\hskip 8.5359pt\textrm{for $\sigma\in 2^{<\omega}$}$
which we call cylinders. Let $\mu$ be the uniform measure on $2^{\omega}$, determined by $\mu(\llbracket\sigma\rrbracket)=2^{-|\sigma|}$. Probability in $2^{\omega}\times 2^{\omega}$ is reduced to $2^{\omega}$ via the measure-preserving $(x,y)\mapsto x\oplus y$. Classes of $\mu$-measure 0 are called null.
A tree is a downward $\preceq$-closed $T\subseteq 2^{<\omega}$ and
  • $T$ is pruned if every $\sigma\in T$ has an extension in $T$
  • $T$ is perfect if every $\sigma\in T$ has two extensions $\tau|\rho$ in $T$.
  • $x$ is a path through $T$ if all of its prefixes belong to $T$.
Let $[T]$ be the class of all paths through $T$.
We use the standard notion of relative computability in terms of Turing machines with oracles from $2^{\omega}$. Turing reducibility $x\leq_{T}z$ means that $x$ is computable from $z$ (is $z$-computable) and is a preorder calibrating $2^{\omega}$ according to computational power in the Turing degrees.
Effectively open sets, also known as $\Sigma^{0}_{1}$ classes, are subsets of $2^{\omega}$ of the form
$\bigcup_{i}\llbracket\sigma_{i}\rrbracket\hskip 8.5359pt\textrm{where $(\sigma_{i})$ is computable.}$.
Effectively closed sets or $\Pi^{0}_{1}$ classes are the complements of $\Sigma^{0}_{1}$ classes. Every $\Pi^{0}_{1}$ class is the set of paths through some computable tree, and vice versa.
If $(\sigma_{n,i})$ is computable then
  • classes $V_{n}=\bigcup_{i}\llbracket\sigma_{n,i}\rrbracket$ are called uniformly $\Sigma^{0}_{1}$
  • $\bigcap_{n}V_{n}$ is called a $\Pi^{0}_{2}$ class and its complement is a $\Sigma^{0}_{2}$ class.
Equivalently, $V\in\Pi^{0}_{2}$ iff there is a computable predicate $P$ such that
$x\in V\iff\forall n\ \exists s\ P(x\upharpoonright_{n},x\upharpoonright_{s}).$
Relativization to an oracle $r$ defines $\Sigma^{0}_{1}(r),\Pi^{0}_{2}(r)$ classes and so on.
A Martin-Löf test is a uniformly $\Sigma^{0}_{1}$ sequence $(V_{n})$ with $\mu(V_{n})\leq 2^{-n}$.
Definition 2.1.
Given $k\in\mathbb{N}$, a real $x$ is
  • random if $x\notin\bigcap_{n}V_{n}$ for any Martin-Löf test $(V_{n})$
  • weakly random if it is in every $\Sigma^{0}_{1}$ set of measure $1$.
  • weakly $k$-random if it is in every $\Sigma^{0}_{k}$ set of measure $1$.
Relativization to oracle $r$ defines $r$-random, weakly $r$-random and so on.
By the countable additivity of the uniform measure $\mu$:
$x$ is weakly $r$-random iff it is not in any null $\Sigma^{0}_{2}(r)$ class (2)
and similarly for weakly $k$-random reals $x$.
2.2 Computable analysis
A Turing machine with a one-way infinite output tape and access to an oracle from $2^{\omega}$ may eventually print an infinite binary sequence, hence defining a partial map from reals to reals. This standard notion of computability of real functions [25] is almost as old as computability itself [10] and implies that computable real functions are continuous. Let
  • $f:\subseteq 2^{\omega}\to 2^{\omega}$ denote that $f$ is a partial map from $2^{\omega}$ to $2^{\omega}$
  • $f(x)\downarrow,f(x)\uparrow$ denote that $f(x)$ is defined or undefined
  • $f(x;n):=f(x)(n)$ denote bit $n$ of $f(x)$
and $f(x;n)\downarrow,f(x;n)\uparrow$ denote that $f(x;n)$ is defined or undefined.
If $f$ is partial computable, the oracle-use of $f(x;n)$ is the prefix of $x$ that has been read by the underlying Turing machine at the time where when $f(x;n)$ is printed on position $n$ of the one-way output tape.
Definition 2.2.
We say that $f:\subseteq 2^{\omega}\to 2^{\omega}$ is random-preserving if $f(x)$ is random for each random $x$ in the domain of $f$.
The range and the inverse images of $f$ are denoted by
$f(2^{\omega}):=\{\hskip 0.85355pt{y}:{\exists x,\ f(x)=y}\hskip 0.85355pt\}\hskip 8.5359pt\textrm{and}\hskip 8.5359ptf^{-1}(y):=\{\hskip 0.85355pt{x}:{f(x)=y}\hskip 0.85355pt\}.$
If $f$ is total and continuous the inverse images
$f^{-1}(\llbracket\sigma\rrbracket):=\{\hskip 0.85355pt{x}:{\sigma\prec f(x)}\hskip 0.85355pt\}$
are closed and effectively closed if $f$ is also computable. Continuous functions $f:\subseteq 2^{\omega}\to 2^{\omega}$ can be defined via a representation: a $\subseteq$-monotone map between cylinders. Formally, let $\lim_{\tau\prec x}\hat{f}(\tau)=z$ denote that
[…]
Given a continuous $f:\subseteq 2^{\omega}\to 2^{\omega}$ we say that $g:\subseteq 2^{\omega}\to 2^{\omega}$:
  • inverts $f$ on $y$ if $g(y)\downarrow$ and $f(g(y))\downarrow=y$
  • is an inversion of $f$ if it is continuous and inverts $f$ on all $y\in f(2^{\omega})$.
We say that $\hat{f}:2^{<\omega}\to 2^{<\omega}$ is a representation of $f$ if
  • $\sigma\preceq\tau\implies\hat{f}(\sigma)\preceq\hat{f}(\tau)$
  • $f(x)\downarrow\iff\lim_{\tau\prec x}\hat{f}(\tau)=f(x)\iff\lim_{\tau\prec x}|{\hat{f}(\tau)}|=\infty$.
Every continuous $f:\subseteq 2^{\omega}\to 2^{\omega}$ has a representation. Every partial computable $f$ has a computable representation. Under an effective coding of the graphs of representations into $2^{\omega}$ we identify them with their codes.
Definition 2.3.
We say that a continuous $f:\subseteq 2^{\omega}\to 2^{\omega}$ computes $z$ and denote it by $z\leq_{T}f$ if every representation of $f$ computes $z$.
Note that by [24] there is no canonical way to assign a representation to each continuous $f$: there are continuous $f$ such that for each representation $z$ of $f$ there is another one of lesser Turing degree.
3 One-way functions
Toward foundational research that is relevant to mathematical practice, Levin proposed a new direction [20]. [^2] This includes 15 drafts developed from 2022 to the end of 2024. He used invertibility of real functions to express the axioms of choice and powerset, and emphasized the significance of its computational hardness in the context of his proposal.
He formally extends the notion of one-way functions to continuous domains and he asks about their existence. One may formalize this notion in several ways which are not equivalent, especially in the case of partial computable maps. In §3.1 we give the formal definition from [20] and explain why it is the correct extension of the discrete notion over weaker alternatives.
In §3.2 we construct a total computable one-way surjection. We do not know whether there is a one-way injection on the reals. In §3.3 we establish properties of one-way real functions relating to this question, setting the basis for the comprehensive analysis in §4.
3.1 Levin’s one-way functions on the reals
In computational complexity a function from strings to strings is one-way if it is easy to compute and hard to invert, even probabilistically.
By Levin [20] a one-way real function $f$ must
  1.    (a)
    be partial computable with domain of positive measure
  2.    (b)
    have no effective probabilistic inversion
  3.    (c)
    satisfy $\mu(f^{-1}(\llbracket\tau\rrbracket))=\textsf{O}\hskip 0.28436pt\big{(}{\mu(\llbracket\tau\rrbracket)}\big{)}$.
Conditions (a), (b) clearly correspond to conditions in the discrete one-way functions. Probabilistic inversions of $f$ are facilitated by $g:\subseteq 2^{\omega}\times 2^{\omega}\to 2^{\omega}$ where the secondary argument represents access to a randomness source. However probability in (b) may refer to the domain or the range of $f$.
Levin [20] chose the latter, interpreting “$g$ inverts $f$” as the event that
$f(g(y,r))\downarrow=y$ on randomly chosen $y\in f(2^{\omega})$, $r\in 2^{\omega}$. (3)
Formally, the probability that $g$ inverts $f$ is the measure of
$L_{f}(g):=\{\hskip 0.85355pt{y\oplus r}:{g(y,r)\downarrow\ \wedge\ \ f(g(y,r))\downarrow=y}\hskip 0.85355pt\}$
and $g$ is a randomized inversion of $f$ if $\mu(L_{f}(g))>0$.
Definition 3.1 (Levin [20]).
We say that $f:\subseteq 2^{\omega}\to 2^{\omega}$ is one-way if it
  • is partial computable, random-preserving with positive domain
  • does not have any partial computable randomized inversion.
The alternative is condition (1) of Gács [9] which, as discussed in §1, is strictly weaker if we assume that $f$ is random-preserving.
The mild measure-preservation condition (c) is equivalent to randomness-preservation and implies $\mu(f(2^{\omega}))>0$ for partial computable $f$ (see [5, §3.2]). To see why (c) is essential consider the weaker conditions:
  1. (i)
    $\mu(f(E))>0$ for each subset $E$ of the domain of $f$ with $\mu(E)>0$
  2. (ii)
    the domain of $f$ contains $R$ with $\mu(R)>0$ and $\mu(f(R))>0$
  3. (iii)
    there is a $\emptyset^{\prime}$-random $x$ such that $f(x)$ is random
and note that (c)$\to$(i)$\to$(ii)$\to$(iii) while none of the converses holds. It is not hard to show (see [5, §3.2]) that each (i), (ii), (iii) is equivalent to (c) up to effective restrictions of $f$. So essentially (c) only asks that $f$ maps at least one sufficiently random real to a random real.
Weaker versions based on combinations of (i), (ii), (iii), $\mu(f(2^{\omega}))>0$ are less robust as they do not form a discernible hierarchy [5].
3.2 Construction of a total one-way surjection
We will map input $x$ to a permutation $y$ of selected bits of $x$. Let
  • $\langle\cdot,\cdot\rangle:\mathbb{N}\times\mathbb{N}\to\mathbb{N}$ be a computable bijection with $\langle n,s\rangle\geq s$
  • $(\emptyset^{\prime}_{s})$ be an effective enumeration of $\emptyset^{\prime}$ without repetitions.
We give a simplified construction which we extend in §3.2 for the full result.
Proposition 3.2.
There exists a total computable $f:2^{\omega}\to 2^{\omega}$ such that any inversion of $f$ computes $\emptyset^{\prime}$.
Proof..
We show that the following $f$ has the required properties:
[…]
Clearly $f$ is total computable. Assuming that $g:\subseteq 2^{\omega}\to 2^{\omega}$ is an inversion of $f$, we show how to decide if $n\in\emptyset^{\prime}$ using the computation of $g(0^{\omega};n)$.
Since $y:=0^{\omega}$ is in the range of $f$, the real $x:=g(y)$ is defined. Let $u_{n}$ be the oracle-use in the computation of $g(y;n)$, that is, the minimum $u$ such that the computation of $g(y;n)$ reads only $y\upharpoonright_{u}$. It remains to show that
  1.    (a)
    if $x(n)=1$ then $n\not\in\emptyset^{\prime}$
  2.    (b)
    if $x(n)=0$ then $n\in\emptyset^{\prime}\iff n\in\emptyset^{\prime}_{u_{n}}$.
If $x(n)=1$ then $\forall s\ f(x;\langle n,s\rangle)=y(\langle n,s\rangle)=0$ so $\forall s\ n\not\in\emptyset^{\prime}_{s}$ and $n\not\in\emptyset^{\prime}$.
For (b) assume that $x(n)=0$ and for a contradiction let $n\in\emptyset^{\prime}_{s}-\emptyset^{\prime}_{s-1}$ for some $s>u_{n}$. Then there is $z\in 2^{\omega}$ with $z=f(0^{n}10^{\omega})=0^{\langle n,s\rangle}10^{\omega}$.
Since $u_{n}$ is the oracle-use of $g(y;n)\downarrow$ and $\langle n,s\rangle\geq s$ we get $0^{u_{n}}\prec z$ and
$g(z;n)=g(0^{\omega};n)=x(n)=0.$
This gives the contradiction $01=z(\langle n,s\rangle)=f(g(z),\langle n,s\rangle)=g(z;n)=$ 0. ∎
Our one-way function will be a permutation of selected bits of the input. We show that such maps meet certain properties required of one-way functions.
Lemma 3.3.
If $p:\mathbb{N}\to\mathbb{N}$ is a computable injection then $f:2^{\omega}\to 2^{\omega}$ with $f(x;n):=x(p(n))$ is a total computable random-preserving surjection.
Proof..
Clearly $f$ is total computable. For each $y$ the real
$x(m):=\begin{cases}y(n)&\textrm{if $m=p(n)$}\\ \ \ 0&\rm{otherwise.}\end{cases}$
satisfies $f(x)=y$. So $f$ is surjective. Let $(V_{i})$ be a universal Martin-Löf test with prefix-free and uniformly c.e. members $V_{i}\subseteq 2^{<\omega}$. Then
$f^{-1}(\llbracket V_{i}\rrbracket)=\bigcup_{\tau\in V_{i}}f^{-1}(\llbracket\tau\rrbracket).$
Since $f$ is computable the sets $f^{-1}(\llbracket V_{i}\rrbracket)$ are uniformly $\Sigma^{0}_{1}$. Also
$f^{-1}(\llbracket\tau\rrbracket)=\{\hskip 0.85355pt{x}:{\forall i<|\tau|,\ x(p(i))=\tau(i)}\hskip 0.85355pt\}$
and since $p$ is injective, $\mu(f^{-1}(\llbracket\tau\rrbracket))=2^{-|\tau|}=\mu(\llbracket\tau\rrbracket)$. So
$\mu\big{(}{f^{-1}(\llbracket V_{i}\rrbracket)}\big{)}\leq\sum_{\tau\in V_{i}}\mu\big{(}{f^{-1}(\llbracket\tau\rrbracket)}\big{)}=\sum_{\tau\in V_{i}}\mu\big{(}{\llbracket\tau\rrbracket}\big{)}=\mu(V_{i}).$
and $(f^{-1}(\llbracket V_{i}\rrbracket))$ is a Martin-Löf test. Since the reals $f$-mapping into $V_{i}$ are in $f^{-1}(\llbracket V_{i}\rrbracket)$ and $(V_{i})$ is universal, every real $f$-mapping to a non-random real is non-random. So $f$ is random-preserving. ∎
We extend the argument in Proposition 3.2 to obtain a one-way function.
Theorem 3.4.
There is a total computable random-preserving one-way surjection, such that every randomized inversion of it computes $\emptyset^{\prime}$.
Proof..
Let $f(x;\langle n,s\rangle):=x(p(\langle n,s\rangle))$ where
$p(\langle n,s\rangle):=\begin{cases}\ \ \ \ 2n&\textrm{if $n\in\emptyset^{\prime}_{s}-\emptyset^{\prime}_{s-1}$}\\[3.69899pt] \ 2\langle n,s\rangle+1&\rm{otherwise}\end{cases}$
so $f$ is total computable and
$f(x;\langle n,s\rangle):=\begin{cases}\ \ \ \ x(2n)&\textrm{if $n\in\emptyset^{\prime}_{s}-\emptyset^{\prime}_{s-1}$}\\[3.69899pt] \ x(2\langle n,s\rangle+1)&\rm{otherwise.}\end{cases}$
Since $p$ is a computable injection, Lemma 3.3 implies that $f$ is a total computable random-preserving surjection. Given $g:\subseteq 2^{\omega}\times 2^{\omega}\to 2^{\omega}$ and
[…]
it remains to show that $g$ computes $\emptyset^{\prime}$. Fix $\sigma$ such that
$\mu(L\cap\llbracket\sigma\rrbracket)>\frac{3}{4}\cdot\mu(\llbracket\sigma\rrbracket)$ (4)
which exists by the Lebesgue density theorem [8, Theorem 1.2.3] or by simply taking an open cover $\llbracket V\rrbracket$ of $L$ such that $\mu(\llbracket V\rrbracket-L)<\mu(L)/3$, where $V=\{\sigma_{0},\sigma_{1},\dots\}$ is prefix-free. Then some $\sigma_{i}$ in $V$ must satisfy (4).
We now compute $\emptyset^{\prime}$ using $g$ and the effective enumeration of a set $W$.
To decide if $n\in\emptyset^{\prime}$, we simultaneously compute $g(y,r;2n)$ for all $y,r$:
if $g(y,r;2n)\downarrow$ with oracle-use $u\geq|\sigma|$ enumerate $(y\oplus r)\upharpoonright_{u}$ in $W$.
Then $L\subseteq\llbracket W\rrbracket$ and $W$ is a c.e. prefix-free set with $\forall\tau\in W,\ |\tau|\geq|\sigma|$.
Effectively in $n$ we produce a computable enumeration $(W_{s})$ of $W$ and
  • compute the least $t$ with $\mu(\llbracket W_{t}\rrbracket\cap\llbracket\sigma\rrbracket)>\mu(\llbracket\sigma\rrbracket)/2$
  • compute the length $k$ of the longest string in $W_{t}$
which exist by (4) and $L\subseteq\llbracket W\rrbracket=\bigcup_{t}\llbracket W_{t}\rrbracket$. It remains to show that
$n\in\emptyset^{\prime}\iff n\in\emptyset^{\prime}_{k}.$ (5)
For a contradiction assume that $n\in\emptyset^{\prime}_{s}-\emptyset^{\prime}_{s-1}$ for some $s>k$.
By the definition of $f$ we have $\forall x,\ f(x;\langle n,s\rangle)=x(2n)$ so
$y\oplus r\in L\implies y(\langle n,s\rangle)=f(g(y\oplus r);\langle n,s\rangle)=g(y\oplus r;2n).$
Fix $\tau\in W_{t}$. Since $\langle n,s\rangle\geq s>k\geq|\tau|$ we have
  • $g(y,r;2n)$ outputs the same result $i$ for $y\oplus r\in\llbracket\tau\rrbracket$
  • only half of the $y\oplus r\in\llbracket\tau\rrbracket$ satisfy $y(\langle n,s\rangle)=i$
so half of $y\oplus r\in\llbracket\tau\rrbracket$ must be outside $L$. Formally:
$\mu(\llbracket\tau\rrbracket-L)\geq\mu(\llbracket\tau\rrbracket)/2$
for each $\tau\in W_{t}$. So $\mu(\llbracket\sigma\rrbracket-L)$ is at least
$\mu(\llbracket W_{t}\rrbracket\cap\llbracket\sigma\rrbracket-L)\geq\mu(\llbracket W_{t}\rrbracket\cap\llbracket\sigma\rrbracket)/2>\mu(\llbracket\sigma\rrbracket)/4$
which contradicts (4), completing the proof of (5). So $g$ computes $\emptyset^{\prime}$.
Since no partial computable $g$ computes $\emptyset^{\prime}$, $f$ is one-way. ∎
one-way functions are not probabilistically invertible on any sufficiently random $z$. We quantify the level of randomness required for this fact.
Proposition 3.5.
There is a random-preserving computable $f:2^{\omega}\to 2^{\omega}$ such that for each $z$, $g:\subseteq 2^{\omega}\times 2^{\omega}\to 2^{\omega}$ satisfying one of:
  1. (i)
    $z$ is weakly 2-random and $g$ is partial computable
  2. (ii)
    $z$ is weakly 1-random and $g$ is total computable
the probability that $g$ inverts $f$ on $z$ is 0.
Proof..
Let $f,L$ be as in Theorem 3.4 and define
$B_{q}\ :=\ \{\hskip 0.85355pt{y}:{\mu(\{\hskip 0.85355pt{r}:{y\oplus r\in L}\hskip 0.85355pt\})\geq q}\hskip 0.85355pt\}.$
Then $\mu(L)=\mu(B_{q})=0$ for each $q\geq 0$. Let $z$ be such that:
$\mu(\{\hskip 0.85355pt{r}:{z\oplus r\in L}\hskip 0.85355pt\})\geq q>0$
for a rational $q>0$ so $z\in B_{q}$. If $g$ is partial computable then
$y\oplus r\in L\iff\forall n\ \exists s\ f(g(y,r);n)[s]\downarrow=y(n)$
so $L$ is $\Pi^{0}_{2}$. Let $(L_{n,s})$ be a computable family of clopen sets with
$L=\bigcap_{n}\bigcup_{s}L_{n,s}\hskip 8.5359pt\textrm{and}\hskip 8.5359pt\bigcup_{s}L_{n+1,s}\subseteq\bigcup_{s}L_{n,s}$
and $L_{n,s}\subseteq L_{n,s+1}$. Then $B_{q}$ is a null $\Pi^{0}_{2}$ class as it is definable by
$y\in B_{q}\iff\forall n\ \exists s\ \mu(\{r:\{y\oplus r\in L_{n,s}\})\geq q.$
By (2) it follows that $z$ is not weakly 2-random. If $g$ is total computable, then $L$ is a null $\Pi^{0}_{1}$ class. By (2) it follows that $z$ is not weakly 1-random. ∎
3.3 Properties of one-way functions
The proof that the total computable $f$ of §3.2 is one-way relied on the fact that $f$ is not injective. This is not a coincidence: since
  • for each tree $T$ with $[T]=\{x\}$ we have $x\leq_{T}T$ uniformly in $T$
  • if $f$ is a total computable injection then $f^{-1}(y)$ consists of the unique path through a tree which is uniformly computable in $y$
total computable injections have computable inverses. More generally:
Theorem 3.6.
If $f:2^{\omega}\to 2^{\omega}$ is total computable, then there exists partial computable $g:\subseteq 2^{\omega}\to 2^{\omega}$ which inverts $f$ in $E:=\{\hskip 0.85355pt{y}:{|{f^{-1}(y)}|=1}\hskip 0.85355pt\}$.
Proof..
Let $\hat{f}$ be a computable representation of $f$. For $y\in E$ the tree
$T_{y}=\{\sigma:\hat{f}(\sigma)\prec y\}$
is uniformly computable in $y$: the map $y\mapsto T_{y}$ is computable. Also
  • $[T_{y}]=f^{-1}(y)$ because $f$ is total
  • if $y\in E$ then $f^{-1}(y)$ is a singleton so $T_{y}$ has a unique path.
The path of any tree $T$ with $|{[T]}|=1$ is computable uniformly in $T$.
So we can compute the unique $x$ in $f^{-1}(y)$ from $y\in E$ uniformly in $y$. ∎
Corollary 3.7.
Every total computable injection $f:2^{\omega}\to 2^{\omega}$ has a total computable inversion $g:2^{\omega}\to 2^{\omega}$.
Partial computable injections need not have computable inverses.
Theorem 3.8.
There exists a partial computable injection $f:\subseteq 2^{\omega}\to 2^{\omega}$ such that any inversion of $f$ computes $\emptyset^{\prime}$.
Proof..
Let $f:\subseteq 2^{\omega}\to 2^{\omega}$ be defined by $f(x):=p(x)\oplus q(x)$ where
$p(x;\langle n,s\rangle):=$$\begin{cases}x(n)&\textrm{if $n\in\emptyset^{\prime}_{s}-\emptyset^{\prime}_{s-1}$}\\[2.84544pt] \ \ 0&\textrm{otherwise}\end{cases}$
$q(x;n):=$$\begin{cases}0&\textrm{if $\forall i\leq n,\ \big{(}{x(i)=0\ \vee\ i\in\emptyset^{\prime}}\big{)}$}\\[2.84544pt] \uparrow&\textrm{otherwise.}\end{cases}$
Then $f$ is partial computable and
  • $f(x)\downarrow$ iff $x$ (as a set of natural numbers) is a subset of $\emptyset^{\prime}$
  • if $f(x)\downarrow$, $f(z)\downarrow$, $x\neq z$ then $x,z$ can only differ on positions in $\emptyset^{\prime}$.
In the latter case $p(x)\neq p(z)$ and $f(x)\neq f(z)$, so $f$ is injective.
As in the proof of Proposition 3.2, every inversion $g$ of $f$ computes $\emptyset^{\prime}$. ∎
We now consider the extent to which total computable one-way functions fail to be injective. Our results hold for a weaker type of one-way functions.
Definition 3.9 (Weakly one-way).
Given $f,h:\subseteq 2^{\omega}\to 2^{\omega}$ let
$L_{f}(h):=\{\hskip 0.85355pt{y}:{h(y)\downarrow\ \wedge\ \ f(h(y))\downarrow=y}\hskip 0.85355pt\}.$
A partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$ is weakly one-way if $\mu(f(2^{\omega}))>0$ and $\mu(L_{f}(h))=0$ for each partial computable $h:\subseteq 2^{\omega}\to 2^{\omega}$.
Clearly one-way functions are weakly one-way.
Corollary 3.10.
Every (weakly) one-way total computable $f:2^{\omega}\to 2^{\omega}$ is almost nowhere injective.
Proof..
Given $f$ as in the statement suppose that
[…]
By Theorem 3.6 there is a partial computable function $h$ with domain $E_{f}$ and $\forall y\in E_{f},\ f(h(y))=y$. So $f$ is not weakly one-way. ∎
For the following, note that Theorem 3.6 relativizes to any cylinder $\llbracket\sigma\rrbracket$.
Theorem 3.11.
If $f:2^{\omega}\to 2^{\omega}$ is a total computable (weakly) one-way function then $f^{-1}(y)$ is uncountable for almost every $y\in f(2^{\omega})$.
Proof..
Suppose that $f:2^{\omega}\to 2^{\omega}$ is total computable. Since countable sets of reals are not perfect it suffices to show that if
$\mu(D)>0\hskip 8.5359pt\textrm{where}\hskip 5.69046ptD:=\{\hskip 0.85355pt{y}:{\textrm{$f^{-1}(y)$ is not perfect}}\hskip 0.85355pt\}$
then $f$ is not weakly one-way. For each $\sigma$ let
  • $f_{\sigma}$ be the restriction of $f$ to $\llbracket\sigma\rrbracket$ and set $E_{\sigma}:=\{\hskip 0.85355pt{y}:{|{f_{\sigma}^{-1}(y)}|=1}\hskip 0.85355pt\}$
  • $g_{\sigma}$ be partial computable which inverts $f_{\sigma}$ on $E_{\sigma}$
where the existence of the $g_{\sigma}$ follows from Theorem 3.6.
If $y\in D$ the closed set $f^{-1}(y)$ has an isolated path, so
$\exists\rho,\ y\in E_{\rho}\hskip 8.5359pt\textrm{and}\hskip 8.5359ptD\subseteq\bigcup_{\sigma}E_{\sigma}.$
Since $\mu(D)>0$ there is $\sigma$ with $\mu(E_{\sigma})>0$ and by the choice of $g_{\sigma}$:
$\mu(\{\hskip 0.85355pt{y}:{g_{\sigma}(y)\downarrow\ \wedge\ \ f(g_{\sigma}(y))=y}\hskip 0.85355pt\})>0.$
This shows that $f$ is not weakly one-way. ∎
4 Inversions of nearly injective functions
Since one-way injections are well-studied in computational complexity [16, 12] it is interesting to ask if there are one-way injections $f$. Their existence remains unknown but as discussed in §3.3, they cannot be total.
With this motivation, we examine the extent to which non-injectivity is essential in the arguments of §3. We exhibit total computable random-preserving surjections that are hard to invert and are nearly injective.
Definition 4.1.
We say that $f:2^{\omega}\to 2^{\omega}$ is two-to-one if $\forall y,\ |{f^{-1}(y)}|\leq 2$.
By extending the method of §3, in §4.1 we exhibit a total computable two-to-one random-preserving surjection whose inversions compute $\emptyset^{\prime}$ but is nevertheless almost everywhere effectively invertible. In §4.2 we exhibit a two-to-one total computable random-preserving surjection with no effective probabilistic map that inverts it almost everywhere.
4.1 Blueprint for two-to-one functions
Our maps will be of the form $f(x\oplus z):=h^{z}(x)\oplus z=y\oplus z$ where
  • $h^{z}$ selects positions of $x$-bits used in (i.e. copied into) $h^{z}(x)$
  • all but at most one position are used in $h^{z}(x)$.
Input: $x\oplus z$;   Output: $y$ with $f(x\oplus z)=y\oplus z$. 1:   Initialization: $k:=0$, $s:=0$ 2:  while true do 3:     if $k\in\emptyset^{\prime}_{s}$ or $E_{s}^{z}(k)$ then 4:        $y(s):=x(k)$ 5:        $k:=s+1$ 6:     else 7:        $y(s):=x(s+1)$ 8:     end if 9:     $s:=s+1$ 10:  end while
Figure 1: Definition of $f$ given a $z$-computable predicate $E^{z}_{s}(i)$.
The selection of $x$-bits is facilitated by a movable marker $k$ and depends on a computable predicate $E^{z}_{s}(i)$ as shown in Figure 1. Let $k^{z}_{s}$ be the candidate for the unique unused position at $s$, which corresponds to $k$ in Figure 1.
The update of $k^{z}_{s}$ occurs if
  • either $k^{z}_{s}\in\emptyset^{\prime}_{s}$ which we call a $\emptyset^{\prime}$-permission
  • or $E^{z}_{s}(k^{z}_{s})$ which we call a $z$-permission
in which case candidate $k^{z}_{s}$ is eliminated. Since $k^{z}_{s}$ is non-decreasing:
  1. (i)
    if $\lim_{s}k^{z}_{s}=\infty$ all $x$-positions are used in $h^{z}(x)$
  2. (ii)
    if $\lim_{s}k^{z}_{s}=k_{0}$ all $x$-positions except $k_{0}$ are used in $h^{z}(x)$.
So $|{f^{-1}(h^{z}(x)\oplus z)}|=1$ if (i) holds and $|{f^{-1}(h^{z}(x)\oplus z)}|=2$ if (ii) holds.
We now need a relativization of Lemma 3.3 that applies to this extended form of selective permutation. To this end, we use a fact from [27]:
$x\oplus y$ is random iff $x$ is random and $y$ is $x$-random (6)
also known as van Lambalgen’s theorem.
Lemma 4.2.
Let $p$ be a total Turing functional such that $n\mapsto p^{z}(n)$ is injective for each oracle $z$ and define $f,h^{z}:2^{\omega}\to 2^{\omega}$ by
$h^{z}(x;n):=x(p^{z}(n))\hskip 8.5359pt\textrm{and}\hskip 8.5359ptf(x\oplus z):=h^{z}(x)\oplus z.$
Then $f$ is a total computable random-preserving surjection.
Proof..
Since $p$ is a total Turing functional, $(z,x)\mapsto h^{z}(x)$ and $f$ are computable. By the relativization of Lemma 3.3 to arbitrary oracle $z$ it follows that $g^{z}$ is a $z$-random preserving surjection.
So $f$ is a surjection. If $x\oplus z$ is random, by (6) we get that
[…]
so $h^{z}(x)\oplus z$ is random. Hence $f$ is random-preserving. ∎
We now apply the above framework to prove:
Theorem 4.3.
There is a total computable $f:2^{\omega}\to 2^{\omega}$ such that:
  • $f$ is a two-to-one random-preserving surjection
  • $f$ is almost everywhere effectively invertible
  • every $g:\subseteq 2^{\omega}\to 2^{\omega}$ that inverts $f$ computes $\emptyset^{\prime}$
and the latter holds for the restriction of $f$ in any cylinder $\llbracket\sigma\rrbracket$.
Proof..
We define $f$ as above with predicate $z(\langle i,s\rangle)=1$ in place of $E^{z}_{s}(i)$.
For each $z$ let $k^{z}_{0}=0$ and
$k^{z}_{s+1}:=\begin{cases}s+1&\textrm{if $k^{z}_{s}\in\emptyset^{\prime}_{s}$ or $z(\langle k^{z}_{s},s\rangle)=1$}\\[3.69899pt] \ \ k^{z}_{s}&\textrm{otherwise.}\end{cases}$ (7)
To select the next $x$-bit used in $f(x\oplus z)$ define:
$p^{z}_{s}:=\begin{cases}s+1&\textrm{if $k^{z}_{s+1}=k^{z}_{s}$}\\[3.69899pt] \ \ k^{z}_{s}&\textrm{otherwise.}\end{cases}$
Then $s\mapsto p^{z}_{s}$ is injective for each $z$. By Lemma 4.2 the $f$ given by
$f(x\oplus z):=h^{z}(x)\oplus z\hskip 8.5359pt\textrm{where}\hskip 8.5359pth^{z}(x;s):=x(p^{z}_{s})$
is a total computable random-preserving surjection. Also
$\forall k,\ |{\{\hskip 0.85355pt{i}:{z(\langle k,i\rangle)=1}\hskip 0.85355pt\}}|=\infty\implies\lim_{s}k^{z}_{s}=\infty\implies|{f^{-1}(h^{z}(x)\oplus z)}|=1$
as discussed above, and $f$ is two-to-one. This condition is met for all random $z$, so $f$ is almost everywhere injective. By Theorem 3.6 there is a partial computable almost everywhere inversion of $f$.
Assuming that $g:\subseteq 2^{\omega}\to 2^{\omega}$ inverts $f$ we show that $g\geq_{T}\emptyset^{\prime}$.
To decide if $n\in\emptyset^{\prime}$, we define $z$ so that $k^{z}_{t}$ gets stuck on $n$ unless $n\in\emptyset^{\prime}$:
$z(\langle i,s\rangle):=\begin{cases}0&\text{if }i=n\\ 1&\text{if }i\neq n.\end{cases}$ (8)
Then $k_{t}^{z}$ does not get stuck on any number $\neq n$ in the sense that
  1.    (a)
    $k_{t}^{z}=n$ at $t=n$
  2.    (b)
    $k_{t}^{z}$ is updated at $t>n$ iff $n\in\emptyset^{\prime}_{t}$.
We show $n\in\emptyset^{\prime}$ iff $n\in\emptyset^{\prime}_{\max\{u,s\}}$, where $u$ is the oracle-use of $g(0^{\omega}\oplus z;2n)$.
For a contradiction suppose $n\in\emptyset^{\prime}_{t}-\emptyset^{\prime}_{t-1}$ for $t>\max\{u,s\}$ so
$\forall x,y\ \ \big{(}{f(x\oplus z)=y\oplus z\implies y(t)=x(n)}\big{)}$ (9)
due to (b) and the definition of $f$. Let $y_{0}:=0^{\omega}$ and $y_{1}:=0^{t}10^{\omega}$.
Then for $i=0,1$ we have $f(g(y_{i}\oplus z))=y_{i}\oplus z$ and by (9):
$g(y_{0}\oplus z;2n)=y_{0}(t)=0\hskip 8.5359pt\textrm{and}\hskip 8.5359ptg(y_{1}\oplus z;2n)=y_{1}(t)=1.$ (10)
Since $u$ is the oracle-use of $g(y_{0}\oplus z_{n};2n)$ and $t>u$ we have
$(y_{0}\oplus z)\upharpoonright_{u}\prec y_{1}\oplus z\hskip 8.5359pt\textrm{so}\hskip 8.5359ptg(y_{0}\oplus z;2n)=g(y_{1}\oplus z;2n)$
which contradicts (10). It follows that $n\in\emptyset^{\prime}\iff n\in\emptyset^{\prime}_{\max\{u,s\}}$.
Finally we modify the above argument so that $g\geq_{T}\emptyset^{\prime}$ is obtained from the weaker assumption that $g$ inverts $f$ inside a cylinder $\llbracket\upsilon\oplus\zeta\rrbracket$. Let
  • $z$ be the extension of $\zeta$ given by (8) for $\langle i,s\rangle\geq|\zeta|$
  • $u$ be the oracle-use of $g(\upsilon 0^{\omega}\oplus z;2n)\downarrow$
  • $y_{0}:=\upsilon 0^{\omega}$ and $y_{1}:=\upsilon 0^{t-|\upsilon|}10^{\omega}$
for $n>|\zeta|$ and $t>u$ which are used for deciding if $n\in\emptyset^{\prime}$ as before.
Assuming $n>|\zeta|$ the modified $z$ satisfies (a), (b). So the above argument applies to the modified $y_{0},y_{1},u$ and proves $g\geq_{T}\emptyset^{\prime}$ as required. ∎
4.2 Almost everywhere probabilistic inversions
By Theorem 3.11 every computable two-to-one random-preserving surjection can be effectively inverted with positive probability. This leaves the possibility that a total computable $f$ exists such that:
  1. (i)
    $f:2^{\omega}\to 2^{\omega}$ is two-to-one, random-preserving and surjective
  2. (ii)
    no partial computable $g$ inverts $f$ with probability 1.
We construct $f$ with the above properties within the framework of §4.1, starting with the modifications and additional ideas needed to achieve this.
In §4.1 we relied on the fact that the given candidate $g$ for inverting $f$ was defined on certain specially constructed computable reals $y_{i}\oplus z_{n}$. This may no longer be the case since we can only assume that $g$ is defined on a set of measure 1. We restrict our considerations to sufficiently random reals.
The domain of a partial computable $g$ which is defined almost everywhere is a $\Pi^{0}_{2}$ class of measure 1 and includes all weakly randoms. In general we only have $g\not\geq_{T}\emptyset^{\prime}$ so for some $r\not\geq_{T}\emptyset^{\prime}$ we use weakly $r$-randoms: reals that are members of every $\Sigma^{0}_{1}(r)$ class of measure 1.
Lemma 4.4.
Suppose that $f:2^{\omega}\to 2^{\omega}$ is a computable surjection and $g:\subseteq 2^{\omega}\to 2^{\omega}$ is an almost everywhere inversion of $f$. If $g\leq_{T}r$ then $g$ inverts $f$ on each weakly $r$-random real.
Proof..
The set of reals where $g$ inverts $f$ is the $\Pi^{0}_{2}(r)$ class
$L_{f}(g):=\{\hskip 0.85355pt{y}:{g(y)\downarrow\ \wedge\ \ f(g(y))=y}\hskip 0.85355pt\}$
which has measure 1 according to the hypothesis. Since weakly $r$-random reals belong to every $\Sigma^{0}_{1}(r)$ class of measure 1, they also belong to every $\Pi^{0}_{2}(r)$ class of measure 1. So $L_{f}(g)$ contains every weakly $r$-random real. ∎
At this point we need some facts about generic reals.
Genericity is a topological form of typicality which has been extensively studied along with randomness in computability [17, 4] and computational complexity [29, 28, 1, 23]. Generic reals avoid every definable meager (as in Baire category) set, as algorithmically random reals avoid definable null set. The level of definability determines the strength of genericity or randomness.
Definition 4.5 (Jockusch [17]).
Given $r,w\in 2^{\omega}$, if
$\exists\sigma\prec w:\ \big{(}{\llbracket\sigma\rrbracket\subseteq G\ \vee\ \llbracket\sigma\rrbracket\cap G=\emptyset}\big{)}.$
for every $\Sigma^{0}_{1}(r)$ class $G\subseteq 2^{\omega}$ we say that $w$ is $r$-generic.
By [17] and [18] (see [8, Theorem 8.11.7]) respectively, for each $r$:
  1.    (a)
    if $r\not\geq_{T}\emptyset^{\prime}$ and $z$ is $r$-generic then $z\oplus r\not\geq_{T}\emptyset^{\prime}$
  2.    (b)
    every $r$-generic is weakly $r$-random and not random.
Definition 4.6.
The $n$th column of $w$ is the real $w^{n}$ with $w^{n}(i):=w(\langle n,i\rangle)$.
By the analogue of (6) for genericity [30, Proposition 2.2] we have
if $w$ is $r$-generic then $w^{n}$ is $r$-generic for each $n$. (11)
By [3, Corollary 2.1] one implication of (6) holds for weak randomness:
if $w$ is weakly $r$-random and $y$ is weakly $w\oplus r$-random then $y\oplus w$ is weakly $r$-random. (12)
Recall that incomputable sets cannot be computed probabilistically with non-zero probability [7]. A relativization of this fact is
$y\not\geq_{T}r\implies\mu\big{(}{\{\hskip 0.85355pt{x}:{x\oplus y\geq_{T}r}\hskip 0.85355pt\}}\big{)}=0$ (13)
which can be found in [8, Corollary 8.12.2].
Given $r\geq_{T}g$, we adapt the argument of §4.1 by choosing $y_{i},z$ so that the $y_{i}\oplus z$ are weakly $r$-random. We construct them from the $y,w$ given by:
Lemma 4.7.
For each $r\not\geq_{T}\emptyset^{\prime}$ there exist $y,w$ such that
  1. (i)
    $y\oplus w$ is weakly $r$-random and $y$ is random
  2. (ii)
    no column $w^{n}$ of $w$ is random
  3. (iii)
    $r\oplus y\oplus w\not\geq_{T}\emptyset^{\prime}$.
Proof..
Let $w$ be $r$-generic and $y$ be $(w\oplus r)^{\prime}$-random so
$y$ is weakly 2-random relative to $w\oplus r$ (14)
and by (b), $w$ is weakly $r$-random. This, combined with (12), gives (i).
By (11) each $w^{n}$ is $r$-generic and by (b) not random, hence (ii).
By (a) we have $w\oplus r\not\geq_{T}\emptyset^{\prime}$ so by (13) the $\Sigma^{0}_{3}(w\oplus r)$ class
[…]
is null. By (14) and (2) we get $y\not\in G$ and (iii) holds. ∎
Fix $w,y$ as in Lemma 4.7 and let $(U_{s})$ be an effective enumeration of a member $U$ of a universal Martin-Löf test with $y\not\in U$ and $\forall n,\ w^{n}\in U$.
For each $n$ we $(y\oplus w)$-effectively define a real $z$ consisting of the columns of $w$ except for the $n$th column which is $y$. Then $y\oplus w$ is weakly $r$-random by (12). Since $z$ copies $y\oplus w$ from a computable array of indices, $z$ is also weakly $r$-random. This suggests defining $E^{z}_{s}(i)$ in the template of §4.1 in terms of the memberships $z^{i}\in U$ of the $i$th column $z^{i}$ of $z$.
A last hurdle in the adaptation of §4.1 is the requirement
from $w\oplus y,n$ compute $s_{n}$ such that $k^{z}_{s_{n}}$ is used iff $n\not\in\emptyset^{\prime}$ (15)
so $k^{z}_{s_{n}}$ does not get $z$-permission. In §4.1 permissions of $k_{s}^{z}$ at $s$ depended entirely on the value of $k_{s}^{z}$ so (15) was achieved by defining $s_{n},z$ with $k_{s_{n}}^{z}=n$. This is no longer possible as we do not have control over the stages $s$ where the $w^{i}$ appear in $U$. The solution is to define permissions in terms of
$d^{z}_{s}:=|{\{\hskip 0.85355pt{k^{z}_{t}}:{t\leq s}\hskip 0.85355pt\}}|$
instead of $k_{s}^{z}$. This will allows to define the required $s_{n},z$ for (15).
Parameters. Given $z$ let $k^{z}_{s}$ be the non-decreasing counter with $k^{z}_{0}=0$ and
$k^{z}_{s+1}:=\begin{cases}s+1&\textrm{if $d^{z}_{s}\in\emptyset^{\prime}_{s}$ or $z^{d^{z}_{s}}\in U_{s}$}\\[3.69899pt] \ \ k^{z}_{s}&\textrm{otherwise}\end{cases}$
[…]
$p^{z}_{s}:=\begin{cases}s+1&\textrm{if $k^{z}_{s+1}=k^{z}_{s}$}\\[3.69899pt] \ \ k^{z}_{s}&\textrm{otherwise}\end{cases}$
enumerates $\mathbb{N}$, omitting $k^{z}_{s}$ as long as it is not updated and including it otherwise. Updates of $k^{z}$ coincide with those of $d^{z}$ and are due to one of:
  • $d^{z}_{s}\in\emptyset^{\prime}_{s}$ which we call $\emptyset^{\prime}$-permission of $k^{z}_{s}$ at $s+1$
  • $z^{d^{z}_{s}}\in U_{s}$ which we call $z$-permission of $k^{z}_{s}$ at $s+1$.
We say that $k^{z}_{s}$ receives permission at $s+1$ if one of the above clauses hold.
Both $d^{z},k^{z}$ are non-decreasing and $d^{z}$ increases by at most 1. The reason that $k^{z}$ is allowed to skip numbers is so that the range of $p^{z}$ misses at most one number $m$, which happens exactly when $\lim_{s}k^{z}_{s}=m$.
So $k^{z}_{s}$ becomes some $p^{z}_{t}$ iff $d^{z}_{s}$ receives permission at a stage $t>s$.
Lemma 4.8.
Given $r\not\geq_{T}\emptyset^{\prime}$ let $y,w$ be as in Lemma 4.7. Effectively in $y\oplus w$ and $n$ we can define $z$ and $s$ such that
  1. (i)
    $d^{z}_{s}=n$ and $\big{(}{\lim_{t}k^{z}_{t}<\infty\iff\lim_{t}k^{z}_{t}=k^{z}_{s}\iff n\not\in\emptyset^{\prime}}\big{)}$
  2. (ii)
    $y\oplus z$ is weakly $r$-random.
Proof..
Let $z$ be the real obtained from $w$ by replacing its $n$th column by $y$. Since $\forall d,\ w^{d}\in U$ each $d\neq n$ will receive $z$-permission at some stage. If $s$ is the stage where each $d<n$ have received permission, $d^{z}_{s}=n$. Since $z^{n}=y\not\in U$ it follows that $n$ will never receive $z$-permission. So $d^{z}$ gets stuck at $n$ if and only if $n$ does not receive $\emptyset^{\prime}$-permission. The required equivalence then follows, given that each $d>n$ receives $z$-permission. ∎
We are now ready to prove:
Theorem 4.9.
There is a total computable $f:2^{\omega}\to 2^{\omega}$ such that:
  • $f$ is a two-to-one random-preserving surjection
  • for each partial $g\not\geq_{T}\emptyset^{\prime}$, with positive probability, $g$ fails to invert $f$
and the latter holds for the restriction of $f$ in any cylinder $\llbracket\sigma\rrbracket$.
Proof..
Define the computable $f:2^{\omega}\to 2^{\omega}$ by
$f(x\oplus z):=h^{z}(x)\oplus z\hskip 8.5359pt\textrm{where}\hskip 8.5359pth^{z}(x;s):=x(p^{z}_{s})$
so $h^{z}(x)$ outputs the bits of $x$ in some order determined by $(\emptyset^{\prime}_{s})$, $z$ with the exception of $\lim_{s}k^{z}_{s}$ when this limit is finite. Then (i), (ii) of §4.1 hold and $f$ is two-to-one. Since $s\mapsto p^{z}_{s}$ is injective, Lemma 4.2 shows that $f$ is a computable random-preserving surjection.
Assuming $g:\subseteq 2^{\omega}\to 2^{\omega}$ inverts $f$ with probability 1, we show that $g\geq_{T}\emptyset^{\prime}$.
For a contradiction assume $g\not\geq_{T}\emptyset^{\prime}$, fix $r\not\geq_{T}\emptyset^{\prime}$ with $g\leq_{T}r$ and
fix $y_{0},v,w$ as in Lemma 4.7 for $y=y_{0}$, so $y_{0}\oplus w\oplus r\not\geq_{T}\emptyset^{\prime}$.
Fix $n$. To decide if $n\in\emptyset^{\prime}$, from $y_{0}\oplus w\oplus r$ we
  • effectively compute $z,s$ as in Lemma 4.8
  • let $k:=k^{z}_{s}$ be the potential finite limit of $k^{z}$
so $y_{0}\oplus z$ is weakly $r$-random, $d^{z}_{s}=n$ and $\forall t>s\ \big{(}{n\in\emptyset^{\prime}_{t}\implies p^{z}_{t}=k}\big{)}$.
By the definition of $f$, for each $x,y$:
$\big{(}{f(x\oplus z)=y\oplus z\ \wedge\ \ t>s\ \wedge\ \ n\in\emptyset^{\prime}_{t}}\big{)}\implies y(t)=x(k).$ (16)
Since $y_{0}\oplus z$ is weakly $r$-random, by Lemma 4.4 we have $g(y_{0}\oplus z)\downarrow$.
Letting $u$ be the oracle-use of $g(y_{0}\oplus z;2k)\downarrow$ we claim that
$n\in\emptyset^{\prime}\iff n\in\emptyset^{\prime}_{\max\{u,s\}}.$ (17)
Otherwise there is $t>\max\{u,s\}$ with $n\in\emptyset^{\prime}_{t}-\emptyset^{\prime}_{t-1}$. Let $y_{1}$ be the real with
$y_{1}(i)=y_{0}(i)\iff i\neq t.$
Then $y_{1}\oplus z$ is weakly $r$-random, so by Lemma 4.4, for $i=0,1$
$g(y_{i}\oplus z)\downarrow\hskip 8.5359pt\textrm{and}\hskip 11.38092ptf(g(y_{i}\oplus z))=y_{i}\oplus z.$ (18)
Since $t>u$ we have $(y_{0}\oplus z)\upharpoonright_{u}\prec y_{1}\oplus z$ so by (17) we get
$g(y_{0}\oplus z;2k)=g(y_{1}\oplus z;2k).$ (19)
By (18), (16) and $t>s$ we get
$g(y_{0}\oplus z;2k)=y_{0}(t)\hskip 8.5359pt\textrm{and}\hskip 8.5359ptg(y_{1}\oplus z;2k)=y_{1}(t)$
which contradict (19) since $y_{0}(t)\neq y_{1}(t)$. This concludes the proof of (17).
By (17) we get $\emptyset^{\prime}\leq_{T}y_{0}\oplus w\oplus r$ which contradicts the hypothesis that $y_{0}\oplus w\oplus r\not\geq_{T}\emptyset^{\prime}$. We conclude that $g\geq_{T}\emptyset^{\prime}$ so every almost everywhere inversion of $f$ computes $\emptyset^{\prime}$. The same argument applies in the case that $g$ inverts $f$ almost everywhere in a cylinder $\llbracket\sigma\rrbracket$. ∎
5 Conclusion
We constructed a one-way function on the reals, which is the analogue of the one-way functions in computational complexity formulated by Levin in [20]. We argued that Levin’s definition is the correct analogue of the one-way functions from computational complexity, over weaker alternatives. An analysis of alternatives was conducted in [5], along with a study of the oracles needed to probabilistically invert one-way real functions.
Our result was based on a general framework for constructing permutations of selected bits of the input, which was adapted toward understanding the extent to which one-way functions can be injective. Applications often require additional ideas, but can yield analogues of significant properties in computational complexity, such as collision-resistance [2].
Despite the non-trivial adaptations that are often required, all currently known one-way real functions (as defined in [20]) are permutations of selected bits of the input and can thus be viewed as applications of our framework. The question therefore arises with respect to the generality of this framework. This is relevant to the existence of a partial computable one-way injection, which is currently unknown.